OpenStudy (anonymous):

Which of the following is equivalent to log{subscript(8)} 24 ? Select one: A. common log of 24 over common log of 8 B. common log of 8 over common log of 24 C. log 3 D. log one-third

OpenStudy (jdoe0001):

\(log_824\) is this what you have?

OpenStudy (anonymous):

\[\log _{a}(b)=\frac{ \log(b) }{ \log(a) }\]\[\log_{8}(24)=?\] @HarvestMan